OpenStudy (anonymous):

sqrt(2/9)... write in form of sqrt(a)/b it says sqrt(2)/3 is incorrect

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{\frac{2}{9}}=\frac{\sqrt{2}}{3}\] is correct, so there has to be a typo somewhere

OpenStudy (anonymous):

there's no typo i shows me what it looks like before i submit the answer, im thinking maybe i'm not suppose to be using square roots is there another way to put it?

jimthompson5910 (jim_thompson5910):

but it says right in the directions to use square roots, so something is wrong maybe they just want the values of 'a' and 'b'?

OpenStudy (radar):

Since the square root is a plus 3 as well as a minus 3 they are asking for\[\sqrt{2}\over \pm3\]

jimthompson5910 (jim_thompson5910):

the "plus/minus" only comes up when solving quadratics

OpenStudy (radar):

Has always been a point of confusion for me.

OpenStudy (radar):

(-3)(-3)=9 The definition of a square root is a factor times itself....

jimthompson5910 (jim_thompson5910):

But the square root is a function. So one input gets mapped to one output. This means that the square of some positive number is positive (and not negative at the same time)

jimthompson5910 (jim_thompson5910):

There's a difference between \[\sqrt{3}\] and \[\pm\sqrt{3}\], the latter results in solving \[x^2=3\]

OpenStudy (radar):

You're right about definition of function.

OpenStudy (anonymous):

i cant enter +/- it wont let me

OpenStudy (radar):

I still can't see why pinksweety gets an incorrect solution!

jimthompson5910 (jim_thompson5910):

yeah that's why I'm thinking it's a typo on their end (ie the program's fault)

OpenStudy (anonymous):

it very well maybe the programs fault, its new this is the first semester they are using it

OpenStudy (radar):

If she had left a radical in the denominator, I can see why they might have docked her for not simplifying putting the radical in the numerator

OpenStudy (radar):

In this case I can't see pinksweety getting a better answer.

jimthompson5910 (jim_thompson5910):

best to just bring it up to an actual human so they can sort this out

OpenStudy (radar):

I would move on pinksweety.

OpenStudy (anonymous):

hang on it says "rationalize the denominator of the expression" does that mean anything?

jimthompson5910 (jim_thompson5910):

the denominator 3 is already a rational number